Find using distributive property: 
5437 × 1001 ????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 5437 × 1001
= 5437 × (1000 + 1)
= 5437 × 1000 + 5437 × 1
= 5437000 + 5437
= 5442437

Thanks